OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x2 - 5x + 5 = 0

OpenStudy (anonymous):

x1=0.7655644371 x2=-3.265564437

OpenStudy (anonymous):

I got -5/4\[-5/4 \pm \sqrt{64} / 4 \] is that right?

OpenStudy (wwe123):

is this correct -2x2 - 5x + 5 = 0 ???

OpenStudy (anonymous):

-2x^2 -5x +5 = 0 Sorry I forgot to put the sqrt in there!

OpenStudy (wwe123):

is √65 and all is right

OpenStudy (anonymous):

So my answer was right just mistaken it for 64 instead 65?

OpenStudy (wwe123):

yes

OpenStudy (wwe123):

√[(-5*-5)-4(-2)(5)] √[25+40] √65

OpenStudy (anonymous):

Ok thanks!